04 March 2013

A sunny start to week two

Week two started with the same measurement-friendly weather we enjoyed last week. The temperature at the Eureka Weather Station remains -47°C; the PEARL Ridge Lab enjoyed the balmy temperature of -25°C.

This morning, Debora and Dan waited at the IR Lab window for the first beams of sunlight to peak over the horizon. Shortly after 9 AM, they were delighted to witness the famous "green flash." Throughout the day at the Ridge Lab, Joseph and Dan acquired 92 MIR measurements with the Bruker, and Debora acquired 51 measurements with PARIS.

Over in the UV-Vis Lab, Xiaoyi analyzed the results of recent UT-GBS tests, which revealed he cannot use the neutral density filter (due to Etalon effect complications). Xiaoyi also finished analysis of the PEARL-GBS' stray light, resolution, polarization, and dark current tests. These tests showed PEARL-GBS is functioning properly, and that the new wedged CCD mount has improved resolution uniformity across the CCD chip. PEARL-GBS is measuring once again in combination mode. SAOZ continues to operate well, and Xiaoyi has sent recent data to Dr. Florence at LATMOS in France. Later in the day, Volodya gave Xiaoyi and Dan a "hands on" lecture about Brewer, which was very interesting.

This morning at the Eureka Weather Station, Volodya, Jim, Emily, and Zen shoveled snow around one of the meteor radar antennae. The snow was 3 feet thick and frozen solid, and care had to be taken not to touch the radar antenna while digging it out. They spent 2 hours clearing snow! Zen, for one, intends to have extra dessert after dinner. Later in the day, Volodya came up to the PEARL Ridge Lab and finished maintenance on Brewer #069 and installed it on the roof with the help from Dan. MAESTRO and SPS are now running on what he hopes will be the final version of the scripts. So far, everything works well.

Inside 0PAL, Zen performed E-AERI system checks. Chris continued to operate the laser all day and all night remotely from Halifax. The CRL group has decided to ship up replacement spare laser pump chambers later this month. There should be no need to use them, provided laser #1 continues to perform nominally, but it is good to have spares on hand. Emily is calibrating depolarization data with the new values measured this year.

Today's ACE ozonesonde flight occurred at 18:15 (local time) and reached an altitude of 32,429 m (6.0 hPa). We have all been very impressed with the clear weather, and hope it continues!

Dan

Back to Daily Reports
Back to Home